Suppose we have a random variable X such that X-1 with probability 1/2 and X =-1 with probability 1/2·We also have another random variable Y such that Y-X with probability 3/4 and YX with probability 1/4. What is the covariance between them, Cov(X, Y)?
Answer #1

Cov(X,Y)- E(XY)-E(XE(Y1) E(X) = 1 × 2-1×2=0 Hence Cov(X, Y)0